# Structure Language Models for Protein Conformation Generation

### v1

@@ 1 Introduction
Protein structure dynamics are fundamental to understanding the biological functions of proteins. The ability of proteins to adopt multiple conformations is crucial for their function in influencing interactions with other biomolecules and the environment. Traditional computational methods, such as molecular dynamics (MD) simulations, have long been used to explore these dynamics. However, these methods are computationally expensive and time-consuming. Structure prediction models, such as AlphaFold 2 (Jumper et al., 2021) and RosettaFold (Baek et al., 2021), have made significant strides in predicting static protein structures, yet often fail to accurately capture the dynamic nature of proteins and their multiple conformations (Chakravarty & Porter, 2022). 000 Equal contribution. Code available at https://github.com/lujiarui/esmdiff. 000  Correspondence to: jiarui.lu@mila.quebec
Recently, significant progress has been made by adopting deep generative models as conformation samplers to efficiently explore the complicated protein conformational space. For example, Noé et al. (2019) adopts normalizing flow to match the underlying Boltzmann distribution by learning from simulation data. Despite their potential, normalizing flow-based methods (Noé et al., 2019; Klein et al., 2023) face significant challenges in modeling large protein systems with hundreds of amino acids, as the invertibility constraint becomes a major obstacle when scaling up model parameters. As a remedy, denoising diffusion approaches (Jing et al., 2023; Lu et al., 2024; Wang et al., 2024; Zheng et al., 2024) can efficiently learn from structural data, achieve good generalization, and perform amortized inference. However, modeling high-dimensional protein structures explicitly in their 3D Euclidean space can demand intensive computation and usually requires accounting for special equivariant properties (Köhler et al., 2020). Furthermore, L2-based training objectives such as denoising score matching (Song et al., 2020) tend to predict local perturbations rather than capturing remote modes of alternative conformations (Wang et al., 2024). Consequently, these models may overallocate their capacity to learn structural noises in the training data instead of focusing on low-frequency structural changes (Chou, 1985).
@@ 2 Related Work
Generative conformation sampling. Given the intensive computation of traditional MD simulations, generative models have been used to learn conformation distributions in a data-driven fashion. The Boltzmann generator (Noé et al., 2019) uses normalizing flow to fit the Boltzmann distribution from target-specific simulation data. Arts et al. (2023) extends this by using denoising diffusion models for coarse-grained protein conformations. Furthermore, EigenFold (Jing et al., 2023), Str2Str (Lu et al., 2024), AlphaFlow (Jing et al., 2024), ConfDiff (Wang et al., 2024), and DiG (Zheng et al., 2024) leverage diffusion or flow matching to conditionally sample protein conformations by learning from PDB data. Recently, AlphaFold3 (Abramson et al., 2024) revised the structure decoder of AlphaFold2 to a diffusion-based module for diversified structure prediction.
@@ 3 Protein Conformation Generation with Language modeling
Notation. A protein with $L$ residues is identified by its sequence of amino acid types ${\bm{c}}\in|{\mathcal{S}}|^{L}$ where ${\mathcal{S}}$ is the vocabulary of 20 standard amino acids. The protein (backbone) structure is represented by its composing 3D atom positions ${\bm{x}}\in{\mathcal{X}}\equiv\mathbb{R}^{L\times 4\times 3}$ including all backbone heavy atoms. Through an encoder $q({\bm{z}}|{\bm{x}})$, the structure ${\bm{x}}$ is encoded to a sequence of latent codes ${\bm{z}}\sim q({\bm{z}}|{\bm{x}})$ where ${\bm{z}}\equiv({\bm{z}}_{1},\dots,{\bm{z}}_{L})\in|V|^{L}$ and $V$ is the pre-specified vocabulary of latent codes; the structure tokens ${\bm{z}}$ are decoded by first mapping to embedding vectors and then to the 3D structure ${\bm{x}}$.
@@ 3.1 Learning the sequence-structure distribution
To address the conformation generation problem, we start with modeling the sequence-to-structure translation distribution $p({\bm{x}}|{\bm{c}})$ of interest and derive the learning objective in this section. To circumvent explicitly learning in the structure space, the roto-translation invariant [^1] For example, features like distance and angle are roto-translation invariant. This relationship can be formally written as $q({\bm{z}}|T\circ{\bm{x}})\triangleq q({\bm{z}}|R\circ{\bm{x}}+t)=q({\bm{z}}|{\bm{x}}),\forall T$. latent representation ${\bm{z}}$ is introduced to encode 3D atomic protein structure. Given this, the target distribution $p({\bm{x}}|{\bm{c}})$ can be derived by marginalizing the joint distribution $p({\bm{x}}|{\bm{c}})=\int_{\bm{z}}p({\bm{x}},{\bm{z}}|{\bm{c}})$. We further factorize this joint distribution according to the Bayes’ rule by isolating the latent variable ${\bm{z}}$: $p_{\theta,\phi}({\bm{x}},{\bm{z}}|{\bm{c}})=p_{\phi}({\bm{x}}|{\bm{c}},{\bm{z}})p_{\theta}({\bm{z}}|{\bm{c}})$, where $p_{\phi}$ denotes the (decoding) distribution over the 3D protein structures given the structure token and sequence, and $p_{\theta}$ denotes the conditional distribution over the structure tokens, respectively modeled by neural networks with parameter set $\phi,\theta$. This gives rise to the evidence lower bound on the likelihood of model distribution over protein structures conditioned on sequence:
@@
where $\psi$ is introduced to parameterized the posterior distribution over latent representation ${\bm{z}}$. Please refer to the Appendix LABEL:app:elbo for the full derivation of Eq. 1. Directly optimizing the right-hand side of Eq. 1 can be intractable and difficult since we have unknown posterior $q_{\psi}$. As a result, we adopt an one-step expectation–maximization (EM) approach (Dempster et al., 1977) by first jointly learning $p_{\phi}$ and $q_{\psi}$ with a simple and parameter-free prior distribution $p({\bm{z}}|{\bm{c}})$, followed by optimization on $p_{\theta}$ with the learned $p_{\phi}^{*}$ and $q_{\psi}^{*}$. This yields the overall two-stage and separable training pipeline:
Learning quantized representation for structure. With the prior $p({\bm{z}}|{\bm{c}})$ fixed, we begin by maximizing the ELBO $\mathcal{L}({\phi,\theta})$ with respect to the encoder $\psi$ and decoder $\phi$, using protein structure samples $\mathcal{D}={{\bm{c}},{\bm{x}}}$. In the context of discrete latent spaces, this process is analogous to training a discrete VAE (dVAE) (Van Den Oord et al., 2017) to learn quantized representations for protein structures. Here, the encoder $q_{\psi}({\bm{z}}|{\bm{x}})$ maps structures to latent tokens, while the decoder $p_{\phi}({\bm{x}}|{\bm{z}},{\bm{c}})$ reconstructs structures from these tokens [^2] We assume that ${\bm{x}}$ is conditionally independent of ${\bm{c}}$ given latent variable ${\bm{z}}$.. The prior $p({\bm{z}}|{\bm{c}})$ is fixed to be uniform during this stage.
Learning the prior over latent tokens. In this stage, we fix the learned parameters $\phi^{*}$ and $\psi^{*}$, and train the prior $p_{\theta}$ by maximizing the ELBO: $\arg\max_{\theta}\mathcal{L}({\phi^{*},\theta})$. Since both $\phi^{*}$ and $\psi^{*}$ are fixed, the reconstruction term in the ELBO cancels out, and training reduces to minimizing the KL divergence $D_{\text{KL}}(q_{\psi^{*}}\|p_{\theta})$. This is equivalent to performing maximum likelihood estimation, as $\mathbb{E}_{({\bm{c}},{\bm{x}})\sim\mathcal{D}}\mathbb{E}_{{\bm{z}}\sim q_{\psi}({\bm{z}}|{\bm{x}})}p_{\theta}({\bm{z}}|{\bm{c}})$ with respect to $p_{\theta}$. Given that both ${\bm{z}}$ and ${\bm{c}}$ are categorical variables, this formulation resembles a translation task, allowing $p_{\theta}$ to be parameterized by language models.
@@
The prior learned in the previous stage is now applied to conformation generation, which can be framed as a conditional generative modeling problem for sequence-to-structure (seq2str) translation. Given an input condition ${\bm{c}}\in|{\mathcal{S}}|^{L}$ which determines the molecular topology, the goal is to sample a conformation ensemble from $p({\bm{x}}|{\bm{c}})$. To do this, we first sample a set of latent variables from the prior distribution learned earlier, ${\bm{z}}\sim p_{\theta}({\bm{z}}|{\bm{c}})$, and then decode these latents using the decoder $p_{\phi}({\bm{x}}|{\bm{c}},{\bm{z}})$. The decoder is jointly trained with the encoder $q_{\psi}({\bm{z}}|{\bm{x}})$ in the first stage, ensuring that the sampled latents align with the reconstruction. This framework supports roto-translation invariant inference and is described in Algorithm 1. Next, we illustrate this approach with two straightforward examples of structure language models (SLM): the encoder-decoder and decoder-only architectures.
@@
Inference involves sampling with a left-to-right decoding order, as defined by the autoregressive factorization of both language models. Figure 3 briefly illustrates these two modeling strategies.
@@ 4 ESMDiff: a Masked Diffusion instantiation
Building on the foundation of SLM, we here propose ESMDiff as an instantiation based on discrete diffusion models (Austin et al., 2021). ESMDiff incorporates the inductive bias of seq2str translation and leverages the protein foundation model ESM3 (Hayes et al., 2024) through masked diffusion fine-tuning. The effectively fine-tuning of ESMDiff also exemplifies how a large pretrained BERT-like masked language model can be adapted to acquire additional generative capabilities, making it well-suited for broader downstream tasks such as conformation generation.
4.1 Revisiting Discrete diffusion as distribution interpolation
The discrete diffusion models (Austin et al., 2021; Lou et al., 2023; Sun et al., 2022; Campbell et al., 2022; Zheng et al., 2023) can be generally defined by a sequential process of progressive noisy variables $z_{t}\in V$ from the categorical variable $z_{0}\in V$. Denote the one-hot (row) vector of $z_{t}$ as ${\bm{z}}_{t}\in\{0,1\}^{|V|}$, in the discrete-time case (Austin et al., 2021), the forward marginal probability of ${\bm{z}}_{t}$ at time $t$ has the following form as a composition of Markov kernel defined by ${\bm{Q}}_{t}~{}(t=1,2\dots,T)$:
$q({\bm{z}}_{t}|{\bm{z}}_{0})=\mathrm{Cat}\left({\bm{z}}_{t};{\bm{z}}_{0}\bar{{\bm{Q}}}_{t}\right)\triangleq\mathrm{Cat}\left({\bm{z}}_{t};{\bm{z}}_{0}{\bm{Q}}_{1}\cdot\dots\cdot{\bm{Q}}_{t}\right),$ (2)
where ${\bm{Q}}_{t}$ indicates the transition probability matrix for time $t$ represented by $[{\bm{Q}}_{t}]_{ij}=q(z_{t}=j|z_{t-1}=i)$, and $\mathrm{Cat}(\cdot;{\bm{p}}),{\bm{p}}\in\Delta^{|V|}$ indicates the categorical distribution with probability and $\Delta^{|V|}$ is the $|V|$-simplex. Eq. 2 also induce the form of the marginal distribution for $\forall t>s$ is $q({\bm{z}}_{t}|{\bm{z}}_{s})=\mathrm{Cat}\left({\bm{z}}_{t};{\bm{z}}_{s}\bar{{\bm{Q}}}_{t|s}\right)\triangleq\mathrm{Cat}\left({\bm{z}}_{t};{\bm{z}}_{s}{\bm{Q}}_{s+1}\cdot\dots\cdot{\bm{Q}}_{t}\right)$. Correspondingly, the posterior $q({\bm{z}}_{s}|{\bm{z}}_{t},{\bm{z}}_{0})$ can be obtained by the reverse process (Austin et al., 2021):
$q({\bm{z}}_{s}|{\bm{z}}_{t},{\bm{z}}_{0})=\frac{q({\bm{z}}_{t}|{\bm{z}}_{s},{\bm{z}}_{0})q({\bm{z}}_{s}|{\bm{z}}_{0})}{q({\bm{z}}_{t}|{\bm{z}}_{0})}=\mathrm{Cat}\left({\bm{z}}_{s};\frac{{\bm{z}}_{t}{{\bm{Q}}}_{t|s}^{\top}\odot{\bm{z}}_{0}\bar{{\bm{Q}}}_{s}}{{\bm{z}}_{0}\bar{{\bm{Q}}}_{t}{\bm{z}}_{t}^{\top}}\right),\forall~{}s<t.$ (3)
Both Zhao et al. (2024) and Shi et al. (2024) discuss how the discrete-time diffusion process can be generalized to the time domain $t\in[0,1]$, akin to the diffusion over continuous space (Song et al., 2020), by demonstrating the continuous-time limit as $T\to\infty$. Notably, when the stationary distribution is explicitly specified (denoted as ${\bm{p}}\in\Delta^{|V|}$), we can choose a state-independent transition kernel in the simple form: ${{\bm{Q}}}_{t|s}\triangleq\left[\alpha(s)^{-1}\alpha(t){\bm{I}}+(1-\alpha(s)^{-1}\alpha(t)){\bm{1}}^{\top}{\bm{p}}\right],$ thus simplifying the continuous-time forward marginal to:
$q({\bm{z}}_{t}|{\bm{z}}_{s})=\mathrm{Cat}\left({\bm{z}}_{t};\frac{\alpha(t)}{\alpha(s)}{\bm{z}}_{s}+(1-\frac{\alpha(t)}{\alpha(s)}){\bm{p}}\right),\forall~{}0\leq s<t<1,$ (4)
where $\alpha(t)\in[0,1)$ is a strictly monotone decreasing function with $\alpha_{0}=1$ and $\alpha_{1}\to 0$. The equation above demonstrates that the discrete diffusion, when defined with an explicit stationary distribution, can be viewed as an interpolation between two categorical distributions controlled by $\alpha(t)$. According to Eq. 3, the reverse process of diffusion defined in Eq. 4 takes the following form for the posterior distribution, where $10\leq s<t<$ 1:
$q({\bm{z}}_{s}|{\bm{z}}_{t},{\bm{z}}_{0})=\mathrm{Cat}\left({\bm{z}}_{s};\frac{\left[\mu(t,s){\bm{z}}_{t}+(1-\mu(t,s))\lambda_{{\bm{z}}_{t}}({\bm{p}}){\bm{1}}\right]\odot\left[\alpha(s){\bm{z}}_{0}+(1-\alpha(s)){\bm{p}}\right]}{\alpha(t)\lambda_{{\bm{z}}_{t}}({\bm{z}}_{0})+(1-\alpha(t))\lambda_{{\bm{z}}_{t}}({\bm{p}})}\right),$ (5)
where $\mu(t,s)\triangleq\alpha(s)^{-1}\alpha(t)>0$ and indicator function $\lambda_{{\bm{z}}_{t}}(\cdot)\triangleq\langle{{\bm{z}}_{t}},\cdot\rangle$ for concision. Readers are referred to Appendix LABEL:app:dd for more details on deriving the Eq. 4 and 5.
Unlike open-ended text generation, protein conformation generation is well-defined within discrete diffusion models, as it conditions on the input amino acid sequence, allowing each output token to correspond uniquely to a position in the input and thus enjoy a fixed-length context window [^3] This indicates that ${\bm{c}}_{[i]}$ and ${\bm{z}}_{[i]}$ are aligned at the same position index $i$..

### v2

@@ Abstract
000 Equal contribution. Code available at https://github.com/lujiarui/esmdiff. 000  Correspondence to: jiarui.lu@mila.quebec, jian.tang@hec.ca.
@@ 1 Introduction
Protein structure dynamics are fundamental to understanding the biological functions of proteins. The ability of proteins to adopt multiple conformations is crucial for their function in influencing interactions with other biomolecules and the environment. Traditional computational methods, such as molecular dynamics (MD) simulations, have long been used to explore these dynamics. However, these methods are computationally expensive and time-consuming. Structure prediction models, such as AlphaFold 2 (Jumper et al., 2021) and RosettaFold (Baek et al., 2021), have made significant strides in predicting static protein structures, yet often fail to accurately capture the dynamic nature of proteins and their multiple conformations (Chakravarty & Porter, 2022).
Recently, significant progress has been made by adopting generative models to efficiently explore the complicated protein conformational space. For example, Noé et al. (2019) adopts normalizing flow to match the underlying Boltzmann distribution by learning from simulation data. Despite their potential, normalizing flow-based methods (Noé et al., 2019; Klein et al., 2023) face challenges in modeling large protein systems with hundreds of amino acids, as the invertibility constraint becomes a major obstacle when scaling up model parameters. As a remedy, denoising diffusion can efficiently learn from structural data (Jing et al., 2023; Lu et al., 2024b; Wang et al., 2024; Zheng et al., 2024), achieve good generalization, and perform amortized inference. However, modeling high-dimensional protein structures explicitly in their 3D Euclidean space can demand intensive computation and usually requires accounting for special equivariant properties (Köhler et al., 2020). Furthermore, L2-based training objectives such as denoising score matching (Song et al., 2020) tend to predict local perturbations rather than capturing remote modes of alternative conformations (Wang et al., 2024). Consequently, these models may overallocate their capacity to learn structural noises in the training data instead of focusing on low-frequency structural changes (Chou, 1985).
@@ 2 Related Work
Generative conformation sampling. Given the intensive computation of traditional MD simulations, generative models have been used to learn conformation distributions in a data-driven fashion. The Boltzmann generator (Noé et al., 2019) uses normalizing flow to fit the Boltzmann distribution from target-specific simulation data. Arts et al. (2023) extends this by using denoising diffusion models for coarse-grained protein conformations. Furthermore, EigenFold (Jing et al., 2023), Str2Str (Lu et al., 2024b), AlphaFlow (Jing et al., 2024a), ConfDiff (Wang et al., 2024), and DiG (Zheng et al., 2024) leverage diffusion or flow matching to conditionally sample protein conformations by learning from PDB data. Recently, AlphaFold3 (Abramson et al., 2024) revised the structure decoder of AlphaFold2 to a diffusion-based module for diversified structure prediction.
@@ 3 Protein Conformation Generation with Language modeling
Notation. A protein with $L$ residues is identified by its sequence of amino acid types ${\bm{c}}\in|{\mathcal{S}}|^{L}$ where ${\mathcal{S}}$ is the vocabulary of 20 standard amino acids. The protein (backbone) structure is represented by its composing 3D atom positions ${\bm{x}}\in{\mathcal{X}}\equiv\mathbb{R}^{L\times 4\times 3}$ including all backbone heavy atoms.
@@ 3.1 Learning the sequence-structure distribution
To address the conformation generation problem, we start with modeling the sequence-to-structure translation distribution $p({\bm{x}}|{\bm{c}})$ of interest and derive the learning objective in this section. To circumvent explicitly learning in the structure space, the roto-translation invariant [^1] For example, features like distance and angle are roto-translation invariant. This relationship can be formally written as $q({\bm{z}}|T\circ{\bm{x}})\triangleq q({\bm{z}}|R\circ{\bm{x}}+t)=q({\bm{z}}|{\bm{x}}),\forall T$. latent representation ${\bm{z}}$ is introduced to encode 3D atomic protein structure, where ${\bm{z}}\equiv({\bm{z}}_{1},\dots,{\bm{z}}_{L})\in|V|^{L}$. Given this, the target distribution $p({\bm{x}}|{\bm{c}})$ can be derived by marginalizing the joint distribution $p({\bm{x}}|{\bm{c}})=\int_{\bm{z}}p({\bm{x}},{\bm{z}}|{\bm{c}})$. We further factorize this joint distribution according to the Bayes’ rule by isolating the latent variable ${\bm{z}}$: $p_{\theta,\phi}({\bm{x}},{\bm{z}}|{\bm{c}})=p_{\phi}({\bm{x}}|{\bm{c}},{\bm{z}})p_{\theta}({\bm{z}}|{\bm{c}})$, where $p_{\phi}$ denotes the (decoding) distribution over the 3D protein structures given the structure token and sequence, and $p_{\theta}$ denotes the conditional distribution over the structure tokens, respectively modeled by neural networks with parameter set $\phi,\theta$. This gives rise to the evidence lower bound on the likelihood of model distribution over protein structures conditioned on sequence:
@@
where $\psi$ is introduced to parameterized the posterior distribution over latent representation ${\bm{z}}$. Please refer to the Appendix LABEL:app:elbo for the full derivation of Eq. (1). Directly optimizing the right-hand side of Eq. (1) can be intractable and difficult since we have unknown posterior $q_{\psi}$. In practice, one may adopt an one-step expectation–maximization (EM) approach (Dempster et al., 1977) by first jointly learning $p_{\phi}$ and $q_{\psi}$ with a simple and parameter-free prior distribution $p({\bm{z}}|{\bm{c}})$, followed by optimization on $p_{\theta}$ with the learned $p_{\phi}^{*}$ and $q_{\psi}^{*}$ similar to Van Den Oord et al. (2017). This yields the overall two-stage and separable training pipeline:
I. Learning quantized representation for structure. With the prior $p({\bm{z}}|{\bm{c}})$ fixed, we begin by maximizing the ELBO $\mathcal{L}({\phi,\theta})$ with respect to the encoder $\psi$ and decoder $\phi$, using protein structure samples $\mathcal{D}=\{({\bm{c}},{\bm{x}})\}$. In the context of discrete latent spaces, this process is analogous to training a discrete VAE (dVAE) (Van Den Oord et al., 2017) to learn quantized representations for protein structures. Here, the encoder $q_{\psi}({\bm{z}}|{\bm{x}})$ maps structures to latent tokens, while the decoder $p_{\phi}({\bm{x}}|{\bm{z}},{\bm{c}})$ reconstructs structures from these tokens [^2] We assume that ${\bm{x}}$ is conditionally independent of ${\bm{c}}$ given latent variable ${\bm{z}}$ for simplicity. In practice, this leads to structure decoder $p_{\phi}({\bm{x}}|{\bm{z}},{\bm{c}})\approx p_{\phi}({\bm{x}}|{\bm{z}})$ such as in Hayes et al. (2024).. The prior $p({\bm{z}}|{\bm{c}})$ is fixed to be uniform during this stage.
II. Learning the prior over latent tokens. In this stage, we fix the learned parameters $\phi^{*}$ and $\psi^{*}$, and train the prior $p_{\theta}$ by maximizing the ELBO: $\arg\max_{\theta}\mathcal{L}({\phi^{*},\theta})$. Since both $\phi^{*}$ and $\psi^{*}$ are fixed, the reconstruction term in the ELBO cancels out, and training reduces to minimizing the KL divergence $D_{\text{KL}}(q_{\psi^{*}}\|p_{\theta})$. This is equivalent to performing maximum likelihood estimation, as $\mathbb{E}_{({\bm{c}},{\bm{x}})\sim\mathcal{D}}\mathbb{E}_{{\bm{z}}\sim q_{\psi}({\bm{z}}|{\bm{x}})}p_{\theta}({\bm{z}}|{\bm{c}})$ with respect to $p_{\theta}$. Given that both ${\bm{z}}$ and ${\bm{c}}$ are categorical variables, this formulation resembles a translation task, allowing $p_{\theta}$ to be parameterized by language models.
Optimizing Eq. (1) provides a general learning framework for conformation generation. In practice, we can approach this objective in a demystified view: we are exploring new “conformations” in an invariant latent space via learning a sequence-to-structure (seq2str) network while offloading the complicated geometric modeling to the structure auto-encoder. This allows the practitioners to choose from popular architectures of structure encoders/decoders and modern language models.
@@
The prior learned in the previous stage is now applied to conformation generation, which can be framed as a conditional generative modeling problem for the seq2str translation. Given an input condition ${\bm{c}}\in|{\mathcal{S}}|^{L}$ which determines the molecular topology, the goal is to sample a conformation ensemble from $p({\bm{x}}|{\bm{c}})$. To do this, we first sample a set of latent variables from the prior distribution learned earlier, ${\bm{z}}\sim p_{\theta}({\bm{z}}|{\bm{c}})$, and then decode these latents using the decoder $p_{\phi}({\bm{x}}|{\bm{c}},{\bm{z}})$. The decoder is jointly trained with the encoder $q_{\psi}({\bm{z}}|{\bm{x}})$ in the first stage, ensuring that the sampled latents align with the reconstruction. This framework supports roto-translation invariant inference and is described in Algorithm 1. Next, we illustrate this approach with two straightforward examples of structure language models (SLM): the encoder-decoder and decoder-only architectures.
@@
Inference involves sampling with a left-to-right decoding order, as defined by the autoregressive factorization of both language models. Fig. 3 briefly illustrates these two modeling strategies.
@@ 4 ESMDiff: a Masked Diffusion instantiation
Building on the foundation of SLM, we here propose ESMDiff as an instantiation based on discrete diffusion models (Austin et al., 2021). ESMDiff incorporates the inductive bias of seq2str translation and leverages the protein foundation model ESM3 (Hayes et al., 2024) through masked diffusion fine-tuning. The effectively fine-tuning of ESMDiff also exemplifies how a large pretrained BERT-like masked language model can be adapted to acquire conditional generative capabilities, making it well-suited for broader downstream tasks such as conformation generation.